# On the interpretation of quantum theory as games between physicists and nature played in Minkowski spacetime

### v1

@@ Abstract
In 2019, we introduced games in Minkowski spacetime as a generalization of game theory to special relativity that subsumes games in normal form (spacelike separation) and games in extensive form (timelike separation). Many concepts including Nash equilibria naturally extend to spacetime games. We also emphasized the importance of these games to model quantum experiments such as Bell experiments and more generally any adaptive measurements. Subsequent work suggested to formalize a special case of such games in terms of strategy presheaves. In the case that measurements have a unique causal bridge and if a natural cover is taken, we show that the two frameworks are isomorphic to each other and provide complementary perspectives. Spacetime games provide a visual and intuitive framework that also captures the distinction between joint experiments and either-or experiments, so that they are rich enough in their causal structure to imply a natural cover for the corresponding causal contextuality scenario. Based on this observation, we suggest to define the strategy presheaf directly based on the pure strategies (and restrictions thereof) of the spacetime game, and we show that the sheaf property obtains for the games at hand. The argument is rather simple and similar to event sheaves for the flat case. Finally, we explain how, in the other direction, the failure of the sheaf property on strategy distribution presheaves is consistent with our previous argument that Nash game theory is not compatible with quantum physics. This shows that the insights of the two frameworks, taken together, can contribute positively to the advancement of the field of quantum foundations.
@@ 1 Introduction and background
Fourny (2019a) and Fourny (2020) introduced a new class of games called spacetime games with perfect information, in order to extend the semantics of game theory to special relativity. In spacetime games, the agents make decisions at various positions in Minkowski spacetime. Spacetime games can be seen as the least common denominator of strategic games on the one hand (spacelike separation), and games in extensive form with perfect information on the other hand (timelike separation).
@@
It is explained in the same reports how these games can be used to model quantum experiments as a game played between experimenters and nature. An example of game corresponding to the Bell experiment is shown in Figure 1 (Alice and Bob are the experimenters and Alfred, which we like to visualize as a cat, represents nature):
Figure 1: A spacetime game with perfect information that represents a Bell experiment with two spacelike experimenters.
@@
A canonical injection of the class of spacetime games with perfect information into the class of games in extensive form with imperfect information is also given. Figure 2 shows the game in extensive form with imperfect information (dashed lines group nodes into information sets, payoffs are omitted) that corresponds to the game of Figure 1. The information sets of the extensive form in Figure 2 correspond exactly to those of the spacetime game in Figure 1 (which are singletons, one for each node, because the spacetime game has perfect information). This is always the case for any spacetime games and their extensive form.
@@
Figure 2: The game in extensive form with imperfect information that correponds to the spacetime shown on Figure 1. Alice and Bob have each one information set for their choice of measurement setting. Alfred has four information sets X, Y, W and Z, one for each measurement.
@@
Subsequent work by Abramsky et al. (2024) suggests to model a subclass of such games, which they call causal contextuality scenarios, as strategy presheaves. In this paper, we show that when causal bridges are unique and when natural covers are taken, causal contextuality scenarios are isomorphic to a subclass of spacetime games, which we call alternating spacetime games. We also show that a natural definition of the strategy presheaf emerges, and that the sheaf property holds. We suggest that the isomorphism can be extended to non-unique causal bridges and to a broader subclass of spacetime games, which we leave open for future work.
1.1 Spacetime games
@@
For the purpose of defining the isomorphism, we now give a compact, formal definition of spacetime games. In this definition, we do not require perfect information. Imperfect information has the meaning proposed by Kuhn (1950) and extends naturally to spacetime games without affecting the injection into the class of games in extensive form with imperfect information. [^1] This is because merging nodes into information sets in the spacetime game translates directly into merging the corresponding information sets in the extensive form.
@@
  • $\mathcal{Z}\subseteq(\mathcal{A}\cup\{\bot\})^{\mathcal{I}}$ is a set of outcomes, which are partial functions from information sets to actions. The outcomes are the complete histories mapping nodes to actions. They must obey some constraints implied by the game structure, and are defined further down in Section 1.3. The domain of definition of an outcome is the set of information sets not mapped to $\bot$, which is the symbol used for unassigned actions.
  • $u=(u_{i})_{i\in P}$ is a family of utility functions mapping outcomes to payoffs, with one payoff for each outcome and player.
The payoffs associated with the outcomes are part of the semantics of a game, but are not the focus of the present paper, which focuses on the structure of the games. They are also not part of any causal contextuality scenario and ignored in the mappings.
Figure 3 shows an example of spacetime game with imperfect information (the dashed lines have the same meaning as on Figure 2) and Figure 4 shows the corresponding extensive form with imperfect information.
Figure 3: A spacetime game with imperfect information that represents a contextuality experiment with two settings to pick out of three. Bob has one information set with three actions (the three contexts XY, XZ and YZ). Alfred has 3 information sets X, Y and Z for each measurement, with two actions each, 0 and 1.
Figure 4: The game in extensive form with imperfect information that corresponds to the spacetime game shown on Figure 3. It has exactly the same information sets: XY, YZ and XZ for Bob, and X, Y and Z for Alfred.
Figure 5 shows an example of spacetime game with perfect information that has a causal bridge and the corresponding extensive form is shown in Figure 6. Spacetime games can be much more complex and mix spacelike-separation as in Figure 1, timelike-separation as in Figure 5, and imperfect information for explicit contexts as shown in Figure 8.
@@
Figure 6: The game in extensive form with imperfect information that corresponds to the spacetime game shown on Figure 5. It happens to have the same tree shape, because there is no spacelike separation
@@ 1.2 Notations
$\textsf{successors}(N)=\{N\in\mathcal{N}|N\smile M\}$
@@
$\textsf{successors}(i)=\bigcup_{N\in i}\textsf{successors}(i)$
@@
We also write $\textsf{parent}(N)$ for the at most unique $P\in\mathcal{N}$ such that $P\smile N$ when it is known from the context that there is a unique predecessor; if it is known that there is none, then this is the empty set. The abuse of notation for information sets then translates to $\textsf{parent}(i)$ where $i$ is an information set for which it is known that there is a unique common predecessor to all its nodes. Finally, when we write the statement $\sigma(N,M)=a$ for some $a$, it implicitly implies $N\smile M$, which allows more concise formulas.
@@ 1.3 Histories
Given a spacetime game in extensive form $\mathcal{G}$, a history is an assignment from $\mathcal{I}$ to $\mathcal{A}\cup\{\bot\}$ if it only assigns an action to activated information sets. An information set is activated if it has a node such that the label of each one of its incoming edges is equal to the action assigned (by that same history) to the information set containing the source node of that edge:
@@
Fourny (2020) uses the term “contingency coordinates” of a node $N$ to denote the minimal history not containing $N$ that coincides with any history in which $N$ is assigned an action (i.e., not $\bot$). In this paper, we use the term “causal bridge” (Fourny et al., 2014) [^2] The idea of a decision happening only if a certain causal past is secured, and otherwise not, can be traced back to (Dupuy, 1992) and (Dupuy, 2000), and we retain his terminology “causal bridge”. of a node $N$ to specifically designates the predecessors of $N$ and the labels on the edges from the predecessors to $N$.
@@
A spacetime game in extensive form $\mathcal{G}$ if valid if the set of outcomes $\mathcal{Z}$ matches the set of complete histories. A complete history is defined as a history that assigns an action to all the activated information sets.
@@
Definition 4 (Clean spacetime game).
A spacetime game is clean if it additionally fulfils the following criteria:
  • It does not contain any information sets that are never assigned any value in any complete history.
    $\forall i\in\mathcal{I},\exists z\in Z,z(i)\neq\bot$
  • It does not contain any unused label.
    $\forall i\in\mathcal{I},\forall a\in\chi(i),\exists z\in Z,z(i)=a$
  • The function $\chi$ does not return unused actions at any information sets, i.e., such that no edge from that node is labeled with them.
    $\forall i\in\mathcal{I},\forall a\in\chi(i),\exists N\in i,\exists M\in\mathcal{N},\sigma(N,M)=a$
  • The set $\mathcal{A}$ does not contain any unused actions.
    $\mathcal{A}=\bigcup_{N\in\mathcal{N}}\chi(N)$
Without loss of generality, we assume the game is clean. It is always possible to remove unused information sets, to remove unused actions from sets returned by $\chi$ and from $\mathcal{A}$, to remove nodes that have an incoming edge with an unused label, and to merge nodes.
@@ 2 Results
We now formally show that, for a subclass of two-player spacetime games (satisfying an “alternating” property) we can build an equivalent causal contextuality scenario. We also show that, in the other direction, we can convert back causal contextuality scenarios to two-player spacetime games. In this paper, we focus on scenarios with unique causal bridges, which means that each measurement is enabled only by one causal past. See Section 3 for a discussion of extensions of the isomorphism (and of the corresponding isomorphic subclasses) to non-unique causal bridges, which are also supported by spacetime games.
Thus, spacetime games and causal contextuality scenarios are isomorphic for a broad class of experimental protocols, which can be either characterized as alternating spacetime games, or as causal contextuality scenarios with unique causal bridges and “good covers” (see Question 7.3 by Abramsky et al. (2024)), whereby we suggest that a cover is “good” is it corresponds to a natural cover implied by a spacetime game structure.
Complete histories under the two frameworks are isomorphic, and histories in causal contextuality scenarios correspond to those histories in spacetime games that have a domain with even cardinality.
@@ 2.1 Alternating spacetime games
We now give a characterization of a subclass of spacetime games called “alternating spacetime games”, for which we will show that there is an isomorphism with the corresponding causal contextuality scenarios.
Note that, since causal contextuality scenarios do not specify which experimenter performs which experiment, we call all our experimenters Bob for the purpose of the mapping. If a causal contextuality scenario is extended with information on the identity of the experimenters (e.g. Alice, Bob) then they can be marked accordingly in the corresponding game.
We include, in the definition of the alternating property, the interpretation of the assumed features of the game in terms of quantum-experimental protocols, to facilitate the understanding of the isomorphism that will be given further down.
Definition 5 (Alternating property).
A spacetime game $\mathcal{G}$ is an alternating game with respect to a player Alfred (typically embodying nature) if it is valid and clean, and:
  1. 1.
    It has two players (we call the other Bob).
  2. 2.
@@
  3. 3.
    All information sets played by Bob are singletons (“unique causal bridge” to the choice of setting).
  4. 4.
    All nodes in the same information set $i$ played by Alfred [^3] This trivially applies to Bob, too. share the same and unique parent node, which we denote $\textsf{parent}(i)$, i.e., all edges to nodes in this information set come from this parent node (“unique causal bridge” to the choice of outcome):
    $\forall i\in\mathcal{I},\exists P\in\mathcal{N},\forall N\in i,\textsf{predecessors}(N)=\{P\}$
  5. 5.
    All information sets played by Bob have at least one successor. In other words, if a context with settings is selected, they the corresponding experiments actually happen and yield a result.
    $\forall N\in\rho^{-1}(\text{Bob}),\textsf{successors}(N)\neq\{\}$
  6. 6.
    $\forall i\in\mathcal{I},\forall N,M\in i,\textsf{successors}(N)=\textsf{successors}(M)$
    $\forall i\in\mathcal{I},\forall N,M\in i,\forall S\in\textsf{successors}(i),\sigma(M,S)=\sigma(N,S)$
  7. 7.
    Two distinct nodes played by Bob cannot have the same causal bridge. [^5] This amounts to saying that we assume that the measurements carried out by Bob under the same enabling circumstances are also “grouped” into a unique experiment where Bob has multiple contexts to choose from. Note that this does not cause any loss of generality, because it is always possible to “fix” a game that would not fulfil this property by merging nodes with the same contingency coordinates, building all combinations of information sets, and concatenating the labels.
    $\forall N,M\in\rho^{-1}(\text{Bob}),$
    $\textsf{predecessors}(N)=\textsf{predecessors}(M)\wedge\forall P\in\textsf{predecessors}(N),\sigma(P,N)=\sigma(P,M)$
    $\implies N=M$
  8. 8.

### v2

# On the interpretation of quantum theory as games between observers and nature played in Minkowski spacetime

@@ Abstract
In 2019, we introduced games in Minkowski spacetime as a generalization of game theory to special relativity that subsumes games in normal form (spacelike separation) and games in extensive form (timelike separation). Many concepts of the game theory literature, including Nash equilibria and flattened strategic forms, naturally extend to spacetime games. We also emphasized the importance of these games to model quantum experiments, such as Bell experiments and more generally any adaptive measurements, from a decision theory perspective. Subsequent work, named causal contextuality scenarios, suggested formalizing a special case of such games in terms of strategy presheaves. If a natural cover is taken, we show that spacetime games recover causal contextuality scenarios as a special case. Spacetime games provide a visual and intuitive framework that captures the distinction between joint experiments and either-or experiments, so that they are rich enough in their causal structure to imply a natural cover for the corresponding causal contextuality scenario. Based on this observation, we suggest defining the strategy presheaf directly based on the pure strategies (and restrictions thereof) in the reduced strategic form of the spacetime game, and we show that the sheaf property obtains for the games at hand. The argument is simple and similar to event sheaves for the flat case. Finally, we explain how, in the other direction, the failure of the sheaf property on strategy distribution presheaves is consistent with our previous argument that Nash game theory is incompatible with quantum physics. This shows that the insights of spacetime games and causal contextuality scenarios, taken together, can contribute positively to the advancement of the field of quantum foundations.
@@ 1.1 Spacetime games and their extensive form
Fourny (2019a) and Fourny (2020) introduced a new class of games called spacetime games with perfect information, to extend the semantics of game theory to special relativity. In spacetime games, the agents make decisions at various positions in Minkowski spacetime. Spacetime games can be seen as the least common denominator of strategic games on the one hand (spacelike separation), and games in extensive form with perfect information on the other hand (timelike separation).
@@
The same reports explain how these games can be used to model quantum experiments as a game played between observers and nature. An example of a game corresponding to the Bell experiment is shown in Figure 1, where Alice and Bob are the observers, and Alfred, whom we like to visualize as a cat, represents nature.
Figure 1: A spacetime game with perfect information that represents a Bell experiment with two spacelike-separated observers. Alice and Bob each select a choice of measurement context (here, only singleton contexts). Alice can pick X or Y, and Bob can pick W or Z. Then, for each one of the two measurements selected, Alfred selects an outcome (0 or 1). Perfect information in spacetime games is suitable for the study of non-locality, a special case of contextuality.
@@
A canonical injection of the class of spacetime games with perfect information into the class of games in extensive form with imperfect information is also given. Figure 2 shows the game in extensive form with imperfect information that corresponds to the game of Figure 1.
1.2 Imperfect information and the dashed lines
Imperfect information is visualized with dashed lines, which group nodes into equivalence classes called information sets. The meaning of these dashes lines is that the agent who makes the decision at these nodes does not know at which one of these nodes they are making their decision: they are not fully informed about the past decisions (in the sense of those made at ancestor nodes in the extensive form), which is why it is called imperfect information. Concretely, in Figure 2, Bob is not informed of Alice’s decision ($\{X\}$ or $\{Y\}$), and this is why his two nodes are connected with a dashed line. This has to do with relativity theory: although Alice may have already made her decision, the information has not reached Bob yet, because his decision is spacelike-separated from Alice’s, as can be seen in the spacelike form of Figure 1, which is basically a causal dependency graph in the sense of relativity theory.
The information sets of the extensive form in Figure 2 correspond exactly to those of the spacetime game in Figure 1 (which are singletons, one for each node, because the spacetime game has perfect information). This is always the case for any spacetime games and their extensive form. The generalization of this canonical injection to all spacetime games, even spacetime games with imperfect information, is straightforward and explained in this paper.
1.3 Imperfect information, locality, and non-contextuality
The semantics of imperfect information and dashed lines has a subtle difference depending on whether one looks at the spacetime game or at the extensive form. In a spacetime game with imperfect information, dashed lines physically represent non-contextuality, connecting identical decisions made in different contexts, whereas in a game in extensive form with imperfect information (in our framework), dashed lines represent either locality or non-contextuality. The spacetime games framework makes explicit that locality is a special case of non-contextuality: in local scenarios, the spacetime game has perfect information and the extensive form has imperfect information, while in more general non-contextualit scenarios, both of them have imperfect information.
@@ 1.3 Imperfect information, locality, and non-contextuality
Figure 2: The game in extensive form with imperfect information that corresponds to the spacetime shown in Figure 1. Alice and Bob have each one information set (A and B) for their choice of measurement setting. Alfred has four information sets X, Y, W, and Z, one for each measurement. For scenarios of interest to the quantum foundations, i.e., exhibiting locality or non-contextuality, the game in extensive form will typically have imperfect information, regardless of whether the spacetime game has perfect information (for locality) or imperfect information (for more general cases of non-contextuality).
Because of the canonical injection of spacetime games into the class of games in extensive form with imperfect information, many well-studied concepts in the game theory literature directly apply to spacetime games. This includes flattening the strategic space with the (reduced) strategic form of a spacetime game (Kuhn, 1950), computing Nash (1951) equilibria, etc.
1.4 Causal contextuality scenarios
@@
Subsequent work by Abramsky et al. (2024) suggests modelling a subclass of such games, which they call causal contextuality scenarios, as strategy presheaves. As we show in this paper, when natural covers are taken, causal contextuality scenarios are isomorphic to a subclass of spacetime games, which we call alternating spacetime games. This means that spacetime games recover causal contextuality scenarios as a special case, and by transitivity, that they also recover the framework of Gogioso and Pinzani as a special case (Gogioso and Pinzani, 2022)(Gogioso and Pinzani, 2023)(Gogioso and Pinzani, 2023), as well as Hardy models (Hardy, 1993), PR boxes (Popescu and Rohrlich, 1994), and GHZ models (Greenberger et al., 1989).
We also show that a flattened definition of the strategy presheaf based on reduced strategic forms (Kuhn, 1950) emerges naturally and that the sheaf property holds.
2 Spacetime games
2.1 Definition
@@
To define the mapping, we now give a formal definition of spacetime games. It is equivalent to that given in (Fourny, 2019a) and (Fourny, 2020), but in a more compact form (see also (Baczyk and Fourny, 2023)). In this definition, we do not require perfect information. Imperfect information has the meaning proposed by Kuhn (1950) and extends naturally to spacetime games without affecting the injection into the class of games in extensive form with imperfect information. [^2] This is because merging nodes into information sets in the spacetime game translates directly into merging the corresponding information sets in the extensive form. Spacetime games with perfect information allow for the study of locality, while spacetime games with imperfect information allow for the more general study of non-contextuality.
@@
  • $\mathcal{Z}\in(\mathcal{A}\cup\{\bot\})^{\mathcal{I}}$ is a set of outcomes, also called complete histories, which are partial functions from information sets to actions. They must obey some constraints implied by the game structure. These constraints are defined further down in Section 2.4. The domain of definition, also called path, of an outcome is the set of information sets not mapped to $\bot$, which is the symbol used for unassigned actions.
  • $u=(u_{i})_{i\in P}$ with, for each $i$, $u_{i}\in\mathbb{R}^{\mathcal{Z}}$ is a family of utility functions mapping outcomes to payoffs, with one payoff for each outcome and player.
The payoffs associated with the outcomes are part of the semantics of a game but are not the focus of the present paper, which focuses on the game structure. They are also not part of any causal contextuality scenario and are ignored in the bijective mapping.
2.2 Further examples
Figure 3 shows an example of a spacetime game with imperfect information (the dashed lines have the same meaning as in Figure 2) and Figure 4 shows the corresponding extensive form with imperfect information.
Figure 3: A spacetime game with imperfect information that represents a non-contextuality experiment with two settings to pick out of three. Bob has one information set with three actions (the three contexts XY, XZ, and YZ). Alfred has 3 information sets X, Y, and Z for each measurement, with two actions each, 0 and 1. This is also known in the literature as a cyclic system of rank 3. Imperfect information in spacetime games captures a more general of non-contextuality than the special case of locality.
Note that, to ease visualization, actions on the leaf nodes are shown as “virtual” edges with no destination node and labeled with the possible actions at these leaf nodes, but these edges are purely visual and do not belong to $\mathcal{R}$; instead, the edges leaving a leaf node $N$ correspond to the elements of $\chi(N)$.
Figure 4: The game in extensive form with imperfect information that corresponds to the spacetime game shown in Figure 3. It has exactly the same information sets: XY, YZ, and XZ for Bob, and X, Y, and Z for Alfred.
Figure 5 shows an example of a spacetime game with perfect information that has a causal bridge and the corresponding extensive form is shown in Figure 6. Spacetime games can be much more complex and mix spacelike-separation as in Figure 1, timelike-separation as in Figure 5, and imperfect information for explicit contexts as shown in Figure 3.
@@ 2.2 Further examples
Figure 6: The game in extensive form with, in this case, perfect information, that corresponds to the spacetime game shown in Figure 5. It happens to have the same tree shape because there is no spacelike separation.
Since locality is a special case of non-contextuality, the Bell experiment can also be represented with a single player picking a context with two measurements, in a cyclic manner (XW, WY, YZ, ZX). The corresponding game is shown in Figure 7. The corresponding game in extensive form is shown in Figure 8 and it is almost the same as the game shown in Figure 2, with the observer players merged into one, which amounts to looking at the Cartesian product of the Hilbert spaces.
Figure 7: A spacetime game with imperfect information that represents a Bell experiment as a non-contextuality experiment instead of a locality experiment, with a single observer player picking a context (A) of two measurement settings out of four possible contexts, and four information sets (X, Y, W, Z) played by nature. This is also known in the literature as a cyclic system of rank 4.
Figure 8: The game in extensive form with imperfect information that corresponds to the spacetime game shown in Figure 7. It has exactly the same information sets: A for Bob, and X, Y, W, and Z for Alfred.
@@ 2.3 Notations
$\textsf{successors}(N)=\{M\in\mathcal{N}|N\smile M\}$
@@
$\textsf{successors}(i)=\bigcup_{N\in i}\textsf{successors}(N)$
@@
When we write the statement $\sigma(N,M)=a$ for some $a$, it implicitly implies $N\smile M$, which allows more concise formulas.
@@ 2.4 Histories
Given a spacetime game in extensive form $\mathcal{G}$, a history is an assignment from $\mathcal{I}$ to $\mathcal{A}\cup\{\bot\}$ that only assigns an action to activated information sets. An information set is activated if it has a node such that the label of each one of its incoming edges is equal to the action assigned (by that same history) to the information set containing the source node of that edge:
@@
Fourny (2020) uses the term “contingency coordinates” of a node $N$ to denote the minimal history not containing $N$ that coincides with any history in which $N$ is assigned an action (i.e., not $\bot$). In this paper, we use the term “causal bridge” (Fourny et al., 2014) [^3] The idea of a decision happening only if a certain causal past is secured, and otherwise not, can be traced back to (Dupuy, 1992) and (Dupuy, 2000), and we retain his terminology “causal bridge”. of a node $N$ to specifically designate the predecessors of $N$ and the labels on the edges from the predecessors to $N$. The two differ in that the contingency coordinates involve the entire ancestry, while the causal bridge only involves (directly connected, one level up) predecessors.
@@
A spacetime game in extensive form $\mathcal{G}$ is valid if the set of outcomes $\mathcal{Z}$ matches the set of complete histories. A complete history is defined as a history that assigns an action to all the activated information sets.
@@
Without loss of expressive power, we assume that games do not contain unused information sets, i.e., there is no information set that never gets activated; otherwise, unused information sets can simply be pruned from the game. [^4] The same assumption will be made for causal contextuality scenarios: No measurement never gets activated, otherwise such measurements can be pruned from the causal contextuality scenario. Likewise, we assume there are no unused actions in $\mathcal{A}$.
@@ 3 Results
We now formally show that, for a subclass of two-player spacetime games (which we call “alternating”), we can build an equivalent causal contextuality scenario. We also show that, in the other direction, we can convert back causal contextuality scenarios to two-player spacetime games.
Thus, spacetime games and causal contextuality scenarios stand in bijection for a broad class of experimental protocols, which can be either characterized as alternating spacetime games or as causal contextuality scenarios with “good covers” (see open question 7.3 by Abramsky et al. (2024)), whereby we suggest that a cover is “good” is it corresponds to the natural cover implied by a spacetime game structure [^5] Abramsky et al. (2024) more recently referred to such good covers as “causally secured covers.”.
Complete histories under the two frameworks are also in bijection, and histories in causal contextuality scenarios correspond to those histories in spacetime games that have a domain with even cardinality (in other words, incomplete histories in causal contextuality scenarios always stop with nature giving an outcome for the chosen measurement, and never end abruptly with a measurement chosen but not carried out).
@@ 3.1 Alternating spacetime games
We now give a characterization of a subclass of spacetime games called “alternating spacetime games”, for which we will show that there is a mapping with the corresponding causal contextuality scenarios.
Note that, since causal contextuality scenarios do not specify which observer performs which experiment, we call all our observers Bob for the mapping. If a causal contextuality scenario is extended with information on the identity of the observers (e.g. Alice, Bob) then they can be marked accordingly in the corresponding game.
We include, in the definition of the alternating property, the interpretation of the assumed features of the game in terms of quantum-experimental protocols, to facilitate the understanding of the mapping that will be given further down.
Definition 4 (Alternating properties).
A spacetime game $\mathcal{G}$ is an alternating game if it is valid and it fulfills the following alternating properties:
  1. 2-PLAYERS
    It has two players (we call them Alfred, who is nature, and Bob, who is the observer).
  2. BIPARTITE
@@
  3. EVEN
    All root nodes are played by Bob. All leaf nodes are played by Alfred.
  4. BOB-S
    All information sets played by Bob are singletons. It means Bob is fully informed about decisions in his causal past, even in situations in which he can carry out the same experiment under different circumstances.
  5. BOB-A
    At any of Bob’s nodes, all available actions are used on at least one edge. [^6] One can also consider allowing for at most one unused action in $\chi(N)$ when looking at causal contextuality scenarios with contexts in which the antichain property does not hold.
    $\forall N\in\rho^{-1}(\text{Bob}),\forall a\in\chi(N),\exists M\in\mathcal{N},\sigma(N,M)=a$
  6. BA1
    Each node played by Alfred has exactly one parent node:
    $\forall N\in\mathcal{N},\rho(N)=\text{Alfred}\implies\exists!P\in\mathcal{N},P\smile N$
  7. BA2